### v1

@@ 3.3.2 Coming-down-from-infinity property
[…]
which goes to 0 as $n\rightarrow\infty$. This proves (41).
The remaining part is devoted to the proof of Proposition 3.34. Let $r=r(n):=\lfloor(\log n)^{1+\varepsilon}\rfloor.$ Fixing $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$, consider the coalescent process
@@
$T^{(r)}_{m}:=\sup\{0\leq j\leq s:X^{(r)}_{j}\leq m\}.$
We begin with a coming-down-from-infinity type estimate.
@@
Assume (H1) and (H2’). Then there exists constant $c>0$ such that for all $n\geq 1$ and $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$,
${\mathbb{P}}_{n}\!\left(T^{(r)}_{\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>s-\frac{n}{\lfloor(\log n)^{1+\frac{\varepsilon}{4}}\rfloor}\right)\;\geq\;1-e^{-c(\log n)^{\frac{\varepsilon}{4}}}.$
Proof 3.38.
By the same argument as in the proof of (42), we have with $\gamma_{m}$ independent $\sim{\rm Geometric}(C^{*}m^{2}/n)$,
${\mathbb{P}}_{n}\!\left(T^{(r)}_{\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\geq\frac{n}{\lfloor(\log n)^{1+\frac{\varepsilon}{4}}\rfloor}\right)\;\leq\;{\mathbb{P}}\!\left(\sum_{m=\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor+1}^{r}\gamma_{m}\;\geq\;\frac{n}{\lfloor(\log n)^{1+\frac{\varepsilon}{4}}\rfloor}\right).$ (51)
Since
${\mathbb{E}}\!\left[\sum_{m=\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor+1}^{r}\gamma_{m}\right]=\sum_{m=\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor+1}^{r}\frac{n}{C^{*}m^{2}}<\frac{n}{C^{*}\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor},$
we obtain
RHS of (51) $\leq{\mathbb{P}}\!\left(\exists\,m=\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor+1,\ldots,r,\,\ \gamma_{m}>C^{*}(\log n)^{\frac{\varepsilon}{4}}\,{\mathbb{E}}[\gamma_{m}]\right)$
$\leq\sum_{m=\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor+1}^{r}{\mathbb{P}}\!\left(\gamma_{m}>C^{*}(\log n)^{\frac{\varepsilon}{4}}\,{\mathbb{E}}[\gamma_{m}]\right)$
$\leq r\,e^{-C^{*}(\log n)^{\frac{\varepsilon}{4}}}\;\leq\;e^{-c(\log n)^{\frac{\varepsilon}{4}}}$
for some constant $c>0$, completing the proof.
Proof 3.39 (Proof of Proposition 3.34).
For $1\leq k\leq(\log n)^{1+\varepsilon}$, the bound (50) follows from Lemma 3.37. We therefore focus on the case $k>(\log n)^{1+\varepsilon}$.
Fix $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$, denote
$j^{*}=j^{*}(s,n):=s-\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}.$
By Theorem 3.14, it suffices to prove that
${\mathbb{P}}_{n}\left(X^{(k)}_{s-{\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}}}\leq\dfrac{k}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\right)\geq 1-e^{-c(\log n)^{\frac{\varepsilon}{4}}}.$ (52)
@@
where $\pi^{n}$ is a uniform permutation in the symmetric group $S_{q_{n}(s)}$. Let $\pi^{n}_{(1)}<\cdots<\pi^{n}_{(r)}$ denote the order statistics of $\{\pi^{n}_{1},\ldots,\pi^{n}_{r}\}$.
Observe that for any $1\leq m\leq r$, the event $\{X^{(r)}_{j^{*}}=m\}$ means that there exist indices
@@
and $m$ distinct vertices $v_{1},\ldots,v_{m}\in\mathcal{X}^{(r)}_{j^{*}}$ such that for each $q=1,\ldots,m$ and every $t\in\{i_{q-1}+1,\ldots,i_{q}\}$,
$\text{Anc}(x^{s}_{\pi^{n}_{(t)}},\,j^{*})=v_{q}.$
That is, the $r$ sampled vertices induce $m$ distinct ancestral lineages at height $j^{*}$, partitioning the index set $\{\pi^{n}_{(1)},\ldots,\pi^{n}_{(r)}\}$ into $m$ consecutive blocks
@@
such that all samples in $I_{q}$ share the same ancestor $v_{q}$ at height $j^{*}$.
On the event $\{X^{(r)}_{j^{*}}=m\}$, this gives a natural decomposition of all vertices at height $s$:
$B_{1}$ $:=\{x^{s}_{i}:\text{Anc}(x^{s}_{i},j^{*})\in\{v_{1},\ldots,v_{m}\}\}=\{x^{s}_{i}:i\in\cup_{q=1}^{m}I_{q}\},$
@@
$\mathbb{P}_{n}\left(X^{(r)}_{j^{*}}\leq\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor\right)\geq{\mathbb{P}}_{n}\!\left(T^{(r)}_{\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>s-\frac{n}{\lfloor(\log n)^{1+\frac{\varepsilon}{4}}\rfloor}\right)\geq 1-e^{-c(\log n)^{\frac{\varepsilon}{4}}}.$ (54)
The next estimate, whose proof will be given later, is the key to completing the argument.
@@
There exists constant $c^{\prime}>0$ such that for any $n\geq 1$, $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$ and $(\log n)^{1+\varepsilon}\leq k\leq q_{n}(s)$,
$\mathbb{P}_{n}\left[\#\big(B_{2}\cap\{u^{s}_{i}:1\leq i\leq k\}\big)>\frac{k}{(\log n)^{\varepsilon/6}}\right]\leq e^{-\frac{c^{\prime}k}{(\log n)^{\varepsilon/6}}}.$ (55)
Substituting (54) and (55) to (53) yields that
$\quad\,{\mathbb{P}}_{n}\left(X^{(k)}_{j^{*}}>\dfrac{k}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\right)$
$\leq\mathbb{P}_{n}\left(X^{(r)}_{j^{*}}>\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor\right)+\mathbb{P}_{n}\left[\#\big(B_{2}\cap\{u^{s}_{i}:1\leq i\leq k\}\big)>\frac{k}{(\log n)^{\varepsilon/6}}\right]$
[…]
for some constant $c^{\prime\prime}>0$. This completes the proof.
Proof 3.41 (Proof of Lemma 3.40).
By (54), it suffices to show that for any $1\leq m\leq\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor$,
${\mathbb{P}}_{n}\!\left(\#\big(B_{2}\cap\{u^{s}_{i}:1\leq i\leq k\}\big)>\frac{k}{(\log n)^{\varepsilon/6}}\,\Big|\,X^{(r)}_{j^{*}}=m\right)\leq e^{-\frac{c^{\prime}k}{(\log n)^{\varepsilon/6}}}.$ (56)
@@
Let $\pi^{n}_{(0)}:=0$ and $\pi^{n}_{(r+1)}:=q_{n}(s)$. Then there exists a constant $c^{\prime}>0$ such that for all $n\geq 1$ and $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$,
${\mathbb{P}}\!\left(\max_{0\leq i\leq r}(\pi^{n}_{(i+1)}-\pi^{n}_{(i)})\leq\frac{n}{(\log n)^{1+\frac{3\varepsilon}{4}}}\right)\geq 1-e^{-c^{\prime}(\log n)^{\frac{\varepsilon}{4}}}.$
On the event $\{X^{(r)}_{j^{*}}=m\}$,
@@
and hence for any $1\leq m\leq\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor$,
${\mathbb{P}}_{n}\!\left(\#B_{2}\leq\frac{n}{(\log n)^{\frac{\varepsilon}{4}}}\,\Big|\,X^{(r)}_{j^{*}}=m\right)\geq 1-e^{-c^{\prime}(\log n)^{\frac{\varepsilon}{4}}}.$ (57)

### v2

@@ 3.3 Coming-down-from-infinity property
$\begin{aligned} &\quad\,{\mathbb{P}}_{n}\left(T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\leq h^{*}-\frac{1}{3}\delta n\right)\leq{\mathbb{P}}_{n}\left(X_{h^{*}-\lfloor\frac{1}{3}\delta n\rfloor}\geq 2(\log n)^{1+\frac{\varepsilon}{2}}\right)\\ &\leq{\mathbb{P}}_{n}\bigg(\#\left\{i=0,\ldots,\left\lfloor\frac{1}{3}\delta(\log n)^{1+\frac{\varepsilon}{8}}\right\rfloor:X_{l_{i+1}}\leq\dfrac{X_{l_{i}}}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\right\}<\frac{1}{6}\delta(\log n)^{1+\frac{\varepsilon}{8}}\bigg).\end{aligned}$ (3.45)
It follows from the Markov property of $X$ in Theorem 3.9 and (3.44) that there exists constant $c>0$ such that for any $n$ sufficiently large and $i=0,\ldots,\left\lfloor\frac{1}{3}\delta(\log n)^{1+\frac{\varepsilon}{8}}\right\rfloor$,
$\mathbb{P}_{n}\left(X_{l_{i+1}}\leq\dfrac{X_{l_{i}}}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\,\Big|\,(X_{j}:l_{i}\leq j\leq h^{*})\right)\geq 1-c(\log n)^{-\frac{13}{24}\varepsilon},\quad\text{a.s..}$
This implies that the RHS of (3.45) goes to 0 as $n\rightarrow\infty$, completing the proof of (3.30). ∎
The remaining part is devoted to the proof of Proposition 3.17. Fix $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{8}}}\leq s\leq h^{*}$. Consider the coalescent process
@@
$T^{(r)}_{m}=T^{(r)}_{m}(s):=\sup\{0\leq j\leq s:X^{(r)}_{j}\leq m\}.$
Denote $r=r(n):=\lfloor(\log n)^{1+\frac{2}{3}\varepsilon}\rfloor$ throughout the proof. For $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{8}}}\leq s\leq h^{*}$, denote
$s^{\downarrow}=s^{\downarrow}(n):=s-\frac{n}{\lfloor(\log n)^{1+\frac{\varepsilon}{8}}\rfloor}.$
We will need the following CDFI-type estimate, whose proof is deferred to the end of this section.
@@
Assume (H1) and (H2a). Then there exists constant $c>0$ such that for all $n$ sufficiently large and $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{8}}}\leq s\leq h^{*}$,
${\mathbb{P}}_{n}\!\left(T^{(r)}_{\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}(s)>s^{\downarrow}\right)\;\geq\;1-c(\log n)^{-\frac{13}{24}\varepsilon}.$
Proof of Proposition 3.17 assuming Lemma 3.19.
For $1\leq k\leq(\log n)^{1+\frac{2}{3}\varepsilon}$, the bound (3.44) follows by Lemma 3.19. We therefore focus on the case $k>(\log n)^{1+\frac{2}{3}\varepsilon}$. By Theorem 3.9, it suffices to prove that
${\mathbb{P}}_{n}\left(X^{(k)}_{s^{\downarrow}}\leq\dfrac{k}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\right)\geq 1-c(\log n)^{-\frac{13}{24}\varepsilon}.$ (3.46)
@@
where $\pi^{n}$ is a uniform permutation in the symmetric group $S_{q_{n}(s)}$. Let $\pi^{n}_{(1)}<\cdots<\pi^{n}_{(r)}$ denote the order statistics of $\{\pi^{n}_{1},\ldots,\pi^{n}_{r}\}$. For any $1\leq m\leq r$, the event $\{X^{(r)}_{s^{\downarrow}}=m\}$ means that there exist indices
@@
and $m$ distinct vertices $v_{1}\prec\cdots\prec v_{m}\in\mathcal{X}^{(r)}_{s^{\downarrow}}$ such that for each $q=1,\ldots,m$ and every $t\in\{i_{q-1}+1,\ldots,i_{q}\}$,
$\text{Anc}(x^{s}_{\pi^{n}_{(t)}},\,s^{\downarrow})=v_{q}.$
That is, the $r$ sampled vertices possess exactly $m$ distinct ancestors, and these ancestral identities partition the index set $\{\pi^{n}_{(1)},\ldots,\pi^{n}_{(r)}\}$ into $m$ consecutive blocks
@@
such that all samples in $I_{q}$ share the same ancestor $v_{q}$ at height $s^{\downarrow}$. On the event $\{X^{(r)}_{s^{\downarrow}}=m\}$, this gives a natural decomposition of all vertices at height $s$:
$B_{1}$ $:=\{x^{s}_{i}:\text{Anc}(x^{s}_{i},s^{\downarrow})\in\{v_{1},\ldots,v_{m}\}\}=\{x^{s}_{i}:i\in\cup_{q=1}^{m}I_{q}\},$
@@
$\mathbb{P}_{n}\left(X^{(r)}_{s^{\downarrow}}\leq\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor\right)\geq{\mathbb{P}}_{n}\!\left(T^{(r)}_{\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>s-\frac{n}{\lfloor(\log n)^{1+\frac{\varepsilon}{8}}\rfloor}\right)\geq 1-e^{-c(\log n)^{\frac{{\varepsilon}}{12}}}.$ (3.48)
To complete the proof, it suffices to prove that
$\mathbb{P}_{n}\left[\#\big(B_{2}\cap\{u^{s}_{i}:1\leq i\leq k\}\big)\leq\frac{k}{(\log n)^{\frac{{\varepsilon}}{6}}}\right]\geq 1-c(\log n)^{-\frac{13}{24}\varepsilon}.$
This follows from (3.48) and the following lemma. ∎
@@
There exists constant $c>0$ such that for any $n$ sufficiently large, $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{8}}}\leq s\leq h^{*}$, $(\log n)^{1+\frac{2}{3}\varepsilon}\leq k\leq q_{n}(s)$ and $1\leq m\leq\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor$,
${\mathbb{P}}_{n}\!\left(\#\big(B_{2}\cap\{u^{s}_{i}:1\leq i\leq k\}\big)>\frac{k}{(\log n)^{\frac{{\varepsilon}}{6}}}\,\Big|\,X^{(r)}_{s^{\downarrow}}=m\right)\leq e^{-c(\log n)^{\frac{{\varepsilon}}{4}}}.$ (3.49)
Proof.
@@
Let $\pi^{n}_{(0)}:=0$ and $\pi^{n}_{(r+1)}:=q_{n}(s)$. Then there exists constant $c^{\prime}>0$ such that for all $n\geq 1$ and $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{8}}}\leq s\leq h^{*}$,
${\mathbb{P}}\!\left(\max_{0\leq i\leq r}(\pi^{n}_{(i+1)}-\pi^{n}_{(i)})\leq\frac{n}{(\log n)^{1+\frac{3}{4}\varepsilon}}\right)\geq 1-e^{-c^{\prime}(\log n)^{\frac{\varepsilon}{4}}}.$
On the event $\{X^{(r)}_{s^{\downarrow}}=m\}$,
@@
and hence by Lemma 3.21, for any $1\leq m\leq\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor$,
${\mathbb{P}}_{n}\!\left(\#B_{2}\leq\frac{n}{(\log n)^{\frac{\varepsilon}{4}}}\,\Big|\,X^{(r)}_{s^{\downarrow}}=m\right)\geq{\mathbb{P}}_{n}\!\left(\#B_{2}\leq\frac{m}{(\log n)^{1+\frac{3}{4}\varepsilon}}\,\Big|\,X^{(r)}_{s^{\downarrow}}=m\right)\geq 1-e^{-c^{\prime}(\log n)^{\frac{\varepsilon}{4}}}.$ (3.50)